OpenStudy (anonymous):

2 5 8 5 7 8 9 17 35 16 1125 ? find the missing term?

OpenStudy (anonymous):

okay, I liked the donkey monkey question ^_^

OpenStudy (anonymous):

7/2=35(3.5x10), 8/5=16(1.6x10), 9/8=1125(1.125*1000), and 17/5=34(3.4*10)

OpenStudy (anonymous):

So 34

OpenStudy (anonymous):

wwow nice logic....

OpenStudy (anonymous):

smart o_o ^_^!